OpenStudy (gabylovesyou):

Let f(x) = 2x + 2. Solve f-1(x) when x = 4.

OpenStudy (anonymous):

that means solve \[2x+2=4\] should easy enough right?

OpenStudy (gabylovesyou):

but inverted...

OpenStudy (gabylovesyou):

f^-1(x)

OpenStudy (anonymous):

yeah i know

OpenStudy (anonymous):

you are not plugging in 4, you are solving for x if \[2x+2=4\]

OpenStudy (anonymous):

takes two steps, subtract 2, divide by 2

OpenStudy (anonymous):

no, when it says to solve for x=2 you insert 2 for x: f(4)=2(4)+2

OpenStudy (gabylovesyou):

x = 1 ?

OpenStudy (anonymous):

trust me on this let me know what you get then i will show you why that answer is right

OpenStudy (anonymous):

er x =4

OpenStudy (gabylovesyou):

i got x = 1 @satellite73

OpenStudy (gabylovesyou):

?

OpenStudy (anonymous):

yes \(x=1\)

OpenStudy (anonymous):

that means that \[f^{-1}(4)=1\]in other words if \(f(x)=2x+2\) then \(f(1)=4\) which is true